# From Slices to Structures: Unsupervised 3D Reconstruction of Female Pelvic Anatomy from Freehand Transvaginal Ultrasound

### v1

@@ 5.3.2 Cross-View Generalization on Real Data
Working with real US data introduces additional challenges, particularly the lack of clear segmentation masks for structures like the uterus. As a result, we evaluate reconstruction quality by testing cross-view generalization. Fig. 12 shows that even when training is done on just one view (sagittal or transversal), the model performs well when tested on the orthogonal view. It also handles difficult regions, such as areas obscured by fluid, with surprising accuracy, which suggests good generalization. Although the model captures anatomical features well, there are alignment issues between sagittal and transversal reconstructions. This complicates training with both views simultaneously. Addressing this misalignment could significantly improve results, as confirmed by the combined-view experiments on synthetic data, where ground truth positions are known. Proper integration of both views may also help estimate slice positions more accurately, potentially moving beyond the simplified assumption of evenly spaced angles. In summary, the method demonstrates strong potential for reconstructing consistent and anatomically meaningful 3D representations from US sweeps.
@@ 5.3.3 Effect of Slice Density
Figure 14: L1 Loss and SSIM across epochs for different numbers of slices, same number of per-slice Gaussians.
@@
5.3.4 Effect of Gaussian Count per Slice
To decouple the effects of slice count and Gaussian density, we fixed the total number of Gaussians ($\approx$55,000) by using an on-slice initialization assuming 200 slices, but trained the model using supervision from only 50 slices. This approach was compared to a baseline where both the initialization and training used 50 slices, resulting in fewer Gaussians ($\approx$13,000). As shown in Fig. 15, the higher-density setup significantly improved reconstruction quality: SSIM increased from 0.866 to 0.9276, and L1 loss decreased from 0.0159 to 0.0111. The trade-off was a 74% increase in training time. These results suggest that increasing the number of Gaussians, even if many are not initialized on the supervised slices, can compensate for sparse supervision, albeit at a higher computational cost.
Figure 15: Slices generated from a sagittal sweep supervised with 50 frames. Left:  13,000 Gaussians; Right:  55,000 Gaussians.
@@ 6 Discussion
Practical Limitations. Our method’s reliance on temporally ordered slices with uniform angular spacing makes it highly effective for smooth, stable acquisitions but also introduces key limitations. The reconstruction quality is sensitive to irregular motion, patient movement, or rapid changes in probe orientation, as these actions disrupt the assumed input structure. Even with CUDA acceleration, the method remains computationally intensive. Near-real-time performance may be achievable with further optimization, but current runtimes and memory demands are high, especially for large numbers of Gaussians or high-resolution data. Finally, while we demonstrated feasibility using both synthetic and real datasets, broader clinical validation is needed. Differences in patient anatomy, imaging conditions, and probe operators were not extensively studied, and generalizing to these variables will be important before the method can be adopted in practice.
Clinical Potential. Despite these limitations, the method shows potential for improving clinical US workflows. During TVS examination the ability to generate accurate 3D reconstructions from freehand sweeps could support better visualization and localization of subtle anatomical abnormalities and enhance surgical planing. Additionally, real-time feedback during scanning could guide clinicians in covering relevant regions more thoroughly, potentially reducing missed findings making scanning less operator dependent.. The synthetic dataset and evaluation pipeline developed as part of this work may also serve as a foundation for future studies, offering reproducible benchmarks for US reconstruction research.
@@ 7 Conclusion
We have introduced a novel, unsupervised framework for reconstructing 3D anatomical volumes from freehand 2D transvaginal US sweeps, without requiring external tracking, ground truth trajectories, or pose supervision. By adapting Gaussian Splatting to the specific physics and geometry of US imaging, and designing a custom differentiable rasterizer, our method enables accurate volumetric reconstruction from sparse, uncalibrated slice data. Through extensive experiments on both synthetic and real-world datasets, we demonstrated that the approach can achieve high-fidelity reconstructions, generalizes well across sweep orientations, and remains robust to various acquisition conditions. We showed that initialization, slice density, and the number of Gaussians are critical factors influencing performance, and proposed effective strategies for optimizing these parameters, including density control and parameter-specific learning rates. While challenges remain—particularly in handling multi-view alignment and achieving real-time performance—the results suggest strong potential for clinical deployment. Our method offers a scalable, hardware-free alternative to conventional 3D US systems and provides a foundation for future research in data-driven, volumetric reconstruction and AI-assisted gynecological imaging.

### v2

@@ 5.3.2 Cross-View Generalization on Real Data
Real US data introduces challenges compared to synthetic environments, particularly the lack of clear segmentation masks. Consequently, we evaluate reconstruction quality through both visual cross-view generalization and quantitative benchmarking against the adapted ImplicitVol baseline.
Fig. 9 demonstrates the qualitative capabilities of our model. Even when training is restricted to a single view (sagittal or transversal), the model generalizes well when tested on the orthogonal view. It captures anatomical features with high fidelity and handles difficult regions, such as areas obscured by fluid, with surprising accuracy.
To strictly quantify this performance, we compared our method against the ImplicitVol baseline on the real clinical data. As shown in Table 5, our approach’s advantages are more pronounced here than in synthetic experiments. Our method outperforms the baseline across all metrics (SSIM, PSNR, LPIPS) for both sagittal and transversal sweeps. Notably, our approach achieves superior perceptual quality (indicated by lower LPIPS) and a drastic reduction in computational cost. While the implicit baseline requires over 3 hours to converge on real data, our method delivers superior results in just 11 to 16 minutes. This order-of-magnitude improvement in efficiency is critical for clinical adoption, where rapid feedback is essential.
Table 5: Quantitative comparison of our method with different number of Gaussians against ImplicitVol on real patient data.
Train Test
SSIM PSNR LPIPS SSIM PSNR LPIPS Time Memory
Sagittal
ImplicitVol 0.858 28.10 0.240 0.783 26.43 0.300 3.14h 10GB
TVGS (Ours) - 50k 0.901 31.38 0.188 0.845 28.51 0.223 11m 129MB
TVGS (Ours) - 100k 0.925 32.24 0.143 0.864 28.85 0.185 16m 160MB
Transversal
ImplicitVol 0.855 27.38 0.247 0.790 26.21 0.295 3.07h 10GB
TVGS (Ours) - 50k 0.897 30.81 0.194 0.847 28.44 0.223 11m 129MB
TVGS (Ours) - 100k 0.923 31.21 0.145 0.823 27.80 0.222 16m 160MB
Despite the strong individual view reconstruction, combining multiple views remains a challenge. We observed alignment discrepancies between sagittal and transversal reconstructions, which complicates training on both views simultaneously. Addressing this misalignment is a key direction for future work; proper integration of both views could further refine slice position estimation, potentially moving beyond the simplified assumption of evenly spaced angles. In summary, our method demonstrates strong potential for reconstructing consistent and anatomically meaningful 3D representations from US sweeps, offering a scalable and fast alternative to existing implicit representations
@@ 5.3.3 Effect of Slice Density
Figure 11: SSIM and LPIPS across epochs for different numbers of slices, same number of per-slice Gaussians.
@@ 6 Discussion
Practical Limitations. Our method relies on initialization from temporally ordered slices with uniform angular spacing. While our experiments on jitter (Table 3) demonstrate that joint pose optimization can effectively compensate for moderate motion irregularities and hand tremor, the method remains sensitive to rapid, large-scale changes in probe orientation that violate the initialization prior. Furthermore, our rasterization framework utilizes a simplified additive model that neglects acoustic attenuation and shadowing. While this trade-off enables faster performance required for operator feedback, it limits the physical realism of the rendering compared to computationally intensive wave-propagation models. Even with CUDA acceleration, the method remains computationally intensive. Near-real-time performance may be achievable with further optimization, but current runtimes are high, especially for large numbers of Gaussians or high-resolution data. Finally, while we demonstrated feasibility using both synthetic and real datasets, broader clinical validation is needed. Differences in patient anatomy, imaging conditions, and probe operators were not extensively studied, and generalizing to these variables will be important before the method can be adopted in practice.
Clinical Potential. Despite these limitations, the method shows potential for improving clinical US workflows. During TVS examination the ability to generate accurate 3D reconstructions from freehand sweeps could support better visualization and localization of subtle anatomical abnormalities and enhance surgical planning. Additionally, real-time feedback during scanning could guide clinicians in covering relevant regions more thoroughly, potentially reducing missed findings making scanning less operator dependent. The synthetic dataset and evaluation pipeline developed as part of this work may also serve as a foundation for future studies, offering reproducible benchmarks for US reconstruction research.
@@ 7 Conclusion
We have introduced a novel, unsupervised framework for reconstructing 3D anatomical volumes from freehand 2D transvaginal US sweeps, without requiring external tracking, ground truth trajectories, or pose supervision. By adapting Gaussian Splatting to the specific physics and geometry of US imaging, and designing a custom differentiable rasterizer, our method enables accurate, fast and memory-efficient volumetric reconstruction from sparse, uncalibrated slice data. Through extensive experiments on both synthetic and real-world datasets, we demonstrated that the approach can achieve high-fidelity reconstructions, generalizes well across sweep orientations, and remains robust to various acquisition conditions. We showed that initialization, slice density, adaptive pose refinement, and the number of Gaussians are critical factors influencing performance, and proposed effective strategies such as density control for optimizing these parameters. By eliminating the reliance on fixed geometric priors, the proposed learnable pose formulation proves that high-fidelity 3D US reconstruction can be achieved as a purely computational, self-correcting task. While challenges remain—particularly regarding multi-view alignment and real-time performance—results suggest strong potential for clinical deployment. Our method offers a scalable, hardware-free alternative to conventional 3D US systems and provides a foundation for future research in data-driven, volumetric reconstruction and AI-assisted gynecological imaging.